OpenStudy (anonymous):

Need help solving these: 1) 3K+10=2K-21 2) 4m-9=5m+7 3) 8y-6=5y+12 4) x-4=6x-19 5) -3p+8=2p-2 6) -t+10=t+4 These are all equations with variables on both sidez. Can any one help me??

OpenStudy (anonymous):

u want ans 4 all ?

OpenStudy (anonymous):

Yes please.

OpenStudy (anonymous):

1. k=-31

OpenStudy (anonymous):

2.m=-16

OpenStudy (anonymous):

1.k=-31 2.m=-16 3.y=6 4.p=2 5.t=3

OpenStudy (anonymous):

4. x=3

OpenStudy (anonymous):

the logic is , try bringing the like terms to one side and solve them. say in the last one 6) -t+10=t+4 , here t is the same variable and the rest all are numbers. so i'll bring numbers to one side and variables to ohter. t+t=10-4 2t=6 so t=3 like wise u atttempt the rest :)

OpenStudy (anonymous):

sry i missed 4.x=3